OpenStudy (anonymous):

Raymond is buying boxes of party favors for his son's birthday party. He has to decide how many boxes he should purchase. A formula for this scenario is c=nz/f where c=number of children at the party n=number of party favors per box z=number of boxes f=number of favors each child will get Rewrite the equation to solve for the number of boxes of party favors Raymond should purchase.

OpenStudy (anonymous):

Correct ? c=nz/f, need to solve for z cf=nz (multiply both sides by f) z=cf/n (divide both sides by n and switch sides)

OpenStudy (anonymous):

@♪Chibiterasu

OpenStudy (31356):

c=nz/f, need to solve for z cf=nz (multiply both sides by f) z=cf/n (divide both sides by n and switch sides) Correct!

OpenStudy (anonymous):

thank you (:

OpenStudy (31356):

Your welcome